OpenStudy (anonymous):

1/8+2/T=17/8T HOW DO I SOLVE FOR "T"

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

first you should try to get both T's to the same side of the equation. so.. 17/8T - 2/T = 1/8

OpenStudy (anonymous):

can some one help me

OpenStudy (anonymous):

then you need to get a common denominator. so if you multiply 2/T by (8/8) then you will get 16/8T

OpenStudy (anonymous):

When you have T's in the denominator -- first find the common denominator and multiply every term by that.

OpenStudy (anonymous):

(8T) 1 (8T)2 (8T) 17 ----- + ----- = ------ 8 T 8T

OpenStudy (anonymous):

Cancel the denominators into the numerators.

OpenStudy (anonymous):

T + 16 = 17

OpenStudy (anonymous):

T = 1

OpenStudy (anonymous):

can u help me plz